# Particle detection performance and Geant4 simulation with low-cost CMOS technology

### v1

@@ 5 Data analysis
The electronics produce a non-negligible noise during the pixel readout, especially when working with low signals. This noise causes, in our case, that 99.48 % of the pixels have a signal detected in the experimental data.
@@
Using the OpenCV (Open Computer Vision) libraries opencv , clusters (i.e., a pixel or groups of pixels with a signal different than zero and adjacent to each other) are searched for in all frames. In our analysis, we will study three observables: the ADC distribution, the cluster size, and the maximum ADC per cluster.
When analyzing the cluster size (defined by its number of pixels), we found that the background noise is mostly present in the smallest clusters of 5 pixels at most, and we cannot differentiate between the signal and the background for clusters of similar size. Therefore, on top of the SNR procedure, we make a cut on the cluster size discarding clusters composed by less than or equal to 3 pixels, removing 97% of the remaining background clusters. Using this filtered sample, we analyze the maximum ADC per cluster. The application of this cut will enable us to compare this reduced data sample, nearly devoid of background clusters, with the Geant4-simulated signal that also underwent the 3-pixel cut.
@@ 6 Results
We present the results of the comparison between our experimental data and the Geant4-simulation of the camera, when it is irradiated by a radioactive source. As we already mentioned, this analysis will be performed for three observables: the ADC distribution, the cluster size, and the maximum ADC per cluster, considering as sources: Sr90 and Cs137. Furthermore, by moving the radioactive sources to various distances, we will investigate the square inverse distance law of irradiation.
In general, the experimental data show the statistical uncertainty, while the simulation uncertainty is the systematic one that comes from the uncertainty in the source activity, which is assumed as 20% error on the source activity, according to the datasheet spectrumtechniques . As we anticipated, the experimental data displayed in all the plots has already been filtered by our SNR procedure.
Fig. 6, compares the ADC distributions at the pixel level for all cluster sizes for the experimental data and its corresponding simulation, for both Sr90 and Cs137. It is important to note that the signal saturation is reached at 1023 ADC, equivalent to 15.48 keV of deposited energy. The background noise only attains up to 115 ADC.
In the plots at the top, the experimental data is presented after applying the SNR procedure. In both plots (left and right), the background is noticeable and located at the first bins. In the case of Sr90, the average number of active pixels for the experimental data, after applying the SNR procedure, is 0.016%, while for the Geant4 signal simulation, it is 0.035%. Analogously, for Cs137, the average number of active pixels for the experimental data is 0.008%, and for the simulation, 0.007%, which is a lower number than for Sr90, even if the activity of the Cs137 source is higher. The electrons of the Sr90 produce more active pixels than the electrons from Cs137 because the former can reach higher energy, being that there are fewer electrons from Cs137 that can deposit energy in the detector in comparison with the electrons coming from Sr90. The latter is clearly illustrated in the plots at the bottom of Figs. 6, in which we are zooming in the region up to 150 ADC, where the ADC distribution for Sr90 is higher than for Cs137.
@@
To remove the background at the first bins, as shown in Figs 6, we cut the events below 100 ADC. After this cut, the background left is minimal, and the Sr90 average number of active pixels is reduced to 0.002%, while for Cs137 reaches 0.0005%, getting for both radioactive sources an excellent agreement in terms of the number of active pixels between the experimental data and the simulation. In Fig. 7 we display the ADC distribution after applying the aforementioned cut. Here, we can see a very good agreement between the ADC distributions from the experimental data and the simulation for both sources. This is confirmed by the reduced chi-square test ($\chi^{2}_{\nu}$) for these ADC distributions, where we obtain a $\chi^{2}_{\nu}$ for Sr90 of 0.55, while for Cs137 is 0.82.
@@
Figure 7: ADC number distribution for the experimental and simulated data with a cut to remove values less than or equal to 100 ADC for Sr90 (left) and Cs 137 (right), the experimental data applying the SNR procedure and comparing with pure simulation at 0 mm between sensor and source, for 500 s exposure time. The vertical dotted line shows the saturation level of the CMOS sensor at 1023 ADC equivalent to 15.48 keV
@@
The top panels in Fig. 8 show the cluster size distribution for experimental data and background after the SNR procedure is applied for both sources. The cluster size distribution shows that the background is present up to the five-pixel cluster, where most of them are located below the three-pixel cluster size. Therefore, on top of the SNR procedure, we apply two cuts: one is for removing clusters with less and equal to a three-pixel cluster size. The other cut removes clusters that have a maximum value of ADC, considering each of its pixels is less than 100 ADC. The application of all these cuts is shown at the bottom panels in Fig. 8. Here, it can be observed the improvement of the agreement between the cluster size distribution for data and simulation after the cut is applied. Considering these distributions, the $\chi^{2}_{\nu}$ value for Sr90 is 2.24, and for Cs137 is 1.26. These values indicate that there is a slight discrepancy between the experimental data and the simulation.
@@
Figure 8: Cluster size distribution for the experimental and simulated data for Sr90 (top left) and Cs 137 (top right), the experimental data applying the SNR procedure and comparing with pure simulation. At the bottom are the same distributions removing the cluster size less than or equal to a three-pixels, and removing clusters that have a maximum value of ADC, considering each of its pixels is less than 100 ADC for Sr90 (bottom left) and Cs 137 (bottom right) at 0mm between sensor and source for 500 s exposure time.
@@
Another parameter we analyze is the distribution of the maximum ADC signal per cluster for the same data set as before, shown in Fig. 9. At the top panels, we have the maximum ADC signal per cluster distribution after applying the SNR procedure. In this case, the simulated data distribution is higher than the experimental data along the entire range of maximum ADC signal per cluster. Also, we appreciate the signal saturation at a maximum of 1023 ADC. At the bottom panels and on top of the SNR procedure, we apply the same two cuts as in the case of the cluster size distributions. As a result, for these maximum ADC signal per cluster distributions, the $\chi^{2}_{\nu}$ value for Sr90 is 2.10, and for Cs137, it is 1.96.
@@
Figure 9: Maximum ADC signal per cluster distribution for the experimental and simulated data for Sr90 (top left) and Cs 137 (top right), the experimental data applying the SNR procedure and comparing with pure simulation. At the bottom are the same distributions removing the cluster size less than or equal to a three-pixels, and removing clusters that have a maximum value of ADC, considering each of its pixels is less than 100 ADC for Sr90 (bottom left) and Cs 137 (bottom right) at 0 mm between sensor and source for 500 s exposure time.
@@
To give the reader a summarized perspective of our results, we present in Table 1, the $\chi^{2}_{\nu}$ for the three observables we study: the ADC distribution, the cluster size, and the maximum ADC per cluster distributions. The SNR filter has been applied to the experimental data for all of these distributions. As mentioned before, an extra cut rejecting pixels with less or equal to 100 ADC units has been applied for the ADC distributions in the data and the simulation. For the cluster size and the maximum ADC per cluster distributions, as mentioned already, we impose two cuts to the simulation and data, these are: one removes cluster size less than or equal to three-pixel size, and the other is to remove clusters that have a maximum value of ADC, considering each of its pixels is less than 100 ADC. As we can see the best agreement we obtained is for the ADC distributions, being very reasonable the ones we got for the cluster size and the maximum ADC per cluster distributions, considering the limitations we face. In that sense, it is crucial to remark that the design characteristics of the OV5647 CMOS are aimed at optimizing a photographic image by artificially amplifying the intensities and other parameters of the pixels. The latter is out of our control since it cannot be handled via standard software.
@@
ADC number (cut >100 ADC) 0.55 0.82
Cluster Size 2.24 1.26
Maximum cluster ADC 2.10 1.96
Table 1: The $\chi^{2}_{\nu}$ tests with 50 bins when comparing the experimental data applying the SNR procedure with simulation. For the ADC distributions with a cut of 100 ADC, and for the distributions related to the number of clusters (i.e., size, and maximum ADC signal) with two cuts to remove the cluster size less than or equal to a three-pixels, and remove clusters that have a maximum value of ADC for Sr90 and Cs137 at 0 mm distance between sensor and source for 500 s exposure time.
@@
Figure 10: Cluster size versus maximum ADC signal per cluster for the experimental and simulated data for Sr90 (left) and Cs 137 (right), the experimental data applying the SNR procedure and comparing with simulation, removing the cluster size less than or equal to a three-pixels, and removing clusters that have a maximum value of ADC, considering each of its pixels is less than 100 ADC at 0 mm between sensor and source for 500 s exposure time.
In Fig. 10, we can observe that the simulated data for these distributions are close to the experimental data for both Sr90 and Cs137. Due to the background noise, the experimental data show more scattered clusters above 30 pixels than in the simulation. The $\chi^{2}_{\nu}$ for these distributions is 1.68 for Sr90 and 1.63 for Cs137, showing a good agreement.
In addition, to verify the inverse square distance law behavior of the Sr90 radioactive source we performed measurements every 2 mm from 0 to 18 mm between the source-detector distance, taking 100 frames per each distance selected. In Fig. 11 we show the average number of clusters at different distances ($z$) for both the experimental data with noise reduction SNR procedure and the simulation and their respective fits to $a/(z+b)^{2}$ (i.e. the inverse square of the distance, where $a$ and $b$ are free parameters)
Figure 11: Average cluster number, as a function of distance ($z$) for the Sr90 source. The experimental data applying the SNR procedure is in red and the pure simulation without cuts is in green. Both data and simulation are fitted to the inverse square function $y=a/(z+b)^{2}$.
@@
To show the expected behavior of the signal with distance, we apply to the experimental data the cut of cluster less than or equal to three pixels to remove the background and no cut to the simulation. If we apply the same cut to the simulation, it would have a larger effect for distances greater than 0 mm, producing an abrupt change, reducing the sample in average to less than 1 cluster after 4 mm.
@@
To better appreciate the behavior with the souce-detector distance we normalize both distributions since the cut to the experimental data lowers the cluster count. Fig. 12 shows that these adjustments follow the inverse square of the distance. The parameters obtained for these fits are $a=4.47\pm 0.27$, $b=2.89\pm 0.11$ for the experimental data, and for the simulation $a=2.99\pm 0.15$, $b=2.20\pm 0.07$, with a $\chi^{2}_{\nu}$ of 0.04 and 0.46 for the fit of experimental and simulated data, respectively. This demonstrates the expected behavior for the source-detector distance for both experimental data and the simulation.
@@
Figure 12: Average cluster number, as a function of distance ($z$) for the Sr90 source. The experimental data applying the SNR procedure and a cut in a cluster less than or equal to three pixels is in red and the pure simulation is in green. Both data are fitted to the inverse square function $y=a/(z+b)^{2}$.
@@
To know if the Sr90 and Cs137 sources can be distinguished, we compare the distributions discussed above at the 0-mm distance. Then, we normalize the histograms to avoid differences in the activities of the sources. Fig. 13, 14, and 15 show the normalized distributions at the pixel and cluster level after applying the cuts. The $\chi^{2}_{\nu}$ for the comparison between the two sources in the experiment and the simulation is shown in Table 2. The average $\chi^{2}_{\nu}$ for the experimental data is 1.52, and for the simulation, it is 2.25. Thus, it is impossible to distinguish between these two radioactive sources for the camera capabilities.
@@
ADC number (cut >100 ADC) 1.46 0.17
Cluster Size 1.79 5.87
Maximum cluster ADC 1.32 0.71
Table 2: The $\chi^{2}_{\nu}$ test with 50 bins when comparing the experimental data applying the SNR procedure, and simulation of Sr90 with Cs137 for the normalized ADC distributions with a cut of 100 ADC and the normalized distributions related to the number of clusters (i.e. size, and maximum ADC signal) with two cuts to remove the cluster size less than or equal to a three-pixels, and remove clusters that have a maximum value of ADC signal per cluster 0 mm distance between source and detector for 500 s exposure time.
@@
Figure 14: Normalized cluster size distribution comparing the sources of Sr90 and Cs137 for the experimental data (left) and the simulation (right), experimental data applying the SNR procedure comparing with simulation, removing the cluster size less than or equal to a three-pixels, and removing clusters that have a maximum value of ADC, considering each of its pixels is less than 100 ADC at 0mm between sensor and source for 500 s exposure time.
Figure 15: Normalized maximum ADC signal per cluster distribution comparing the sources of Sr90 and Cs137 for the experimental data (left) and the simulation (right), experimental data applying the SNR procedure comparing with simulation, removing the cluster size less than or equal to a three-pixels, and removing clusters that have a maximum value of ADC, considering each of its pixels is less than 100 ADC at 0mm between sensor and source for 500 s exposure time

### v2

@@ 5 Data analysis
The electronics of the camera produce a non-negligible noise during the pixel readout, especially when working with low signals. This noise causes, in our case, that 99.48 % of the pixels have a signal detected in the experimental data.
@@
Using the OpenCV (Open Computer Vision) libraries [26], clusters (i.e., a pixel or groups of pixels with a signal different than zero and adjacent to each other) are searched for in all frames. In our analysis, we will study three observables: the ADC distribution, the cluster size, and the maximum ADC per cluster.
When analyzing the cluster size (defined by its number of pixels), we found that the background noise is mostly present in the smallest clusters of 5 pixels at most, and we cannot differentiate between the signal and the background for clusters of similar size. Therefore, on top of the SNR procedure, we make a cut on the cluster size discarding clusters composed by less than or equal to 5 pixels, removing 100% of the remaining background clusters. Using this filtered sample, we analyze the maximum ADC per cluster. The application of this cut will enable us to compare this reduced data sample, nearly devoid of background clusters, with the Geant4-simulated signal that also underwent the 5-pixel cluster cut.
@@ 6 Results
We present the results of the comparison between our experimental data and the Geant4-simulation of the camera when it is irradiated by a radioactive source. As we already mentioned, this analysis will be performed for three observables: the ADC distribution, the cluster size, and the maximum ADC per cluster, considering as sources: Sr90 and Cs137. Furthermore, by moving the radioactive sources to various distances, we will investigate the square inverse distance law of irradiation.
In general, the experimental data show the statistical uncertainty, while the simulation uncertainty propagates the systematic source activity uncertainty, assumed to be a 20% error according to the datasheet [17], combined with the standard deviation from ten simulations. As mentioned before, the experimental data shown in all the plots has already been filtered by our SNR procedure.
Figure 6, compares the ADC distributions at the pixel level for all cluster sizes for the experimental data and its corresponding simulation, for both Sr90 and Cs137. It is important to note that the signal saturation is reached at 1023 ADC, equivalent to 15.48 keV of deposited energy. The background noise only attains up to 115 ADC.
In the plots at the top, the experimental data is presented after applying the SNR procedure. In both plots (left and right), the background is noticeable and located at the first bins. In the case of Sr90, the average number of active pixels for the experimental data, after applying the SNR procedure, is 0.011%, while for the Geant4 signal simulation, it is 0.013%. Analogously, for Cs137, the average number of active pixels for the experimental data is 0.005%, and for the simulation, 0.003%, which is a lower number than for Sr90, even if the activity of the Cs137 source is higher.
The electrons emitted by Sr90 have a higher maximum energy than those from Cs137, resulting in a greater number of detections with the Sr90 source than with the Cs137 source, since the higher energy electrons have a higher probability of passing through the different layers of the material until arriving to the detector. In comparison, Cs137 electrons, having lower energy, have a lower probability of depositing energy in the detector [27, 28]. The latter is clearly illustrated in the plots at the bottom of Figs. 6, in which we are zooming in the region up to 150 ADC, where the ADC distribution for Sr90 is higher than for Cs137.
@@
To remove the background at the first bins after applying the SNR procedure, as shown in Figs 6, we cut the events below 100 ADC, since we cannot tell apart signal from background. After this cut, the background left is minimal, and the Sr90 average number of active pixels is reduced to 0.002%, while for Cs137 reaches 0.0005%, getting for both radioactive sources an excellent agreement in terms of the number of active pixels between the experimental data and the simulation.
In figure 7 we display the ADC distribution after applying the aforementioned cut. Here, we can see a very good agreement between the ADC distributions from the experimental data and the simulation for both sources. This is confirmed by the reduced chi-square test ($\chi^{2}_{\nu}$) for these ADC distributions, where we obtain a $\chi^{2}_{\nu}$ for Sr90 of 0.45, while for Cs137 is 0.48.
@@
Figure 7: ADC number distribution for the experimental and simulated data with a cut to remove values less than or equal to 100 ADC for Sr90 (left) and Cs 137 (right), the experimental data applying the SNR procedure and comparing with pure simulation at 0 mm between sensor and source, for 500 s exposure time. The vertical dotted line shows the saturation level of the CMOS sensor at 1023 ADC equivalent to 15.48 keV.
To analyze the effect of the ADC cut and saturation on the signal, we analyze its dynamic range from simulation, shown in figure 8. The energy deposited by electrons and photons emitted from Sr90 and Cs137 decays covers a much larger spectrum than the range detectable with the sensor (from the lower 100 ADC cut (1.5 keV) to the maximum 1023 ADC (15.48 keV)). In the sensitive range of the sensor the normalized distributions of both radioactive sources have the same shape. Only at the highest energies there is a difference. However that area is not reachable for the sensor, since it lies beyond the saturation peak. The lower energy depositions below the 100 ADC cut follow the same trend for Sr90 and Cs137. Thus, even if we are cutting a large part of the low energy signal where the background overlaps, we are not losing differentiation capabilities. In contrast, the energy distributions from the initial emitted electrons and photons do have a different shape, however, it is not translated into the deposited energy.
Figure 8: Normalized energy distributions, without cuts, of electrons and photons emitted from Sr90 and Cs137 and their corresponding deposited energy in the sensor pixel matrix, from the Geant4 simulation. The vertical dashed lines limit the sensor sensitivity, showing the low ADC cut that removes background and the upper ADC sensor saturation.
@@
The top panels in figure 9 show the cluster size distribution for experimental data and background after the SNR procedure is applied for both sources. The cluster size distribution shows that the background is present up to the five-pixel cluster, where most of them are located below the five-pixel cluster size. Therefore, on top of the SNR procedure, we apply two cuts: one is for removing clusters with less or equal to a five-pixel cluster. The other cut eliminates clusters where one or more of its pixels is less than or equal to 100 ADC. We will call the combination of both cuts Cluster Cuts (CC). The 100 ADC cut will also remove part of the signal and modify the structure and total deposited energy of the original clusters. However, since the sensor has no energy resolution capabilities, as explained later, the drawback of this cut is reduced.
The application of all these cuts is shown in the bottom panels of figure 9. Here, we observe the improvement of the agreement between the cluster size distribution for data and simulation after the cuts are applied. Considering these distributions, the $\chi^{2}_{\nu}$ value for Sr90 is 1.77, and for Cs137 is 1.16. These values indicate that there is a slight discrepancy between the experimental data and the simulation.
@@
Figure 9: Cluster size distribution for the experimental and simulated data for Sr90 (top left) and Cs 137 (top right), the experimental data applying the SNR procedure and comparing with pure simulation. At the bottom are the same distributions applying the CC procedure for Sr90 (bottom left) and Cs 137 (bottom right) at 0mm between sensor and source for 500 s exposure time.
@@
Another parameter we analyze is the distribution of the maximum ADC signal per cluster for the same data set as before, shown in figure 10. In the upper panels, we have the maximum ADC signal per cluster distribution after applying the SNR procedure. In this case, the simulated data distribution shows greater efficiency since it is higher than the experimental data along the entire range of maximum ADC signal per cluster.
In addition, we appreciate the signal saturation at a maximum of 1023 ADC. At the bottom panels we apply the SNR and CC procedures. As a result, the $\chi^{2}_{\nu}$ for the maximum ADC signal per cluster distributions is 2.62 for Sr90 and 1.94 for Cs137.
@@
Figure 10: Maximum ADC signal per cluster distribution for the experimental and simulated data for Sr90 (top left) and Cs 137 (top right), the experimental data applying the SNR procedure and comparing with pure simulation. At the bottom are the same distributions applying the CC procedure for Sr90 (bottom left) and Cs 137 (bottom right) at 0 mm between sensor and source for 500 s exposure time.
@@
To give the reader a summarized perspective of our results, we present in Table 1, the $\chi^{2}_{\nu}$ for the three observables we study: the ADC distribution, the cluster size, and the maximum ADC per cluster distributions. The SNR filter has been applied to the experimental data for all of these distributions. As mentioned before, an extra cut rejecting pixels with less than or equal to 100 ADC units has been applied for the ADC distributions in the data and the simulation. For the cluster size and the maximum ADC per cluster distributions, as mentioned already, we apply the CC procedure for simulation and data.
As we can see, the best agreement we obtained is for the ADC distributions, being very reasonable the ones we got for the cluster size and the maximum ADC per cluster distributions, considering the limitations we face. In that sense, it is crucial to remark that the design characteristics of the OV5647 CMOS are aimed at optimizing a photographic image by artificially amplifying the intensities and other parameters of the pixels. The latter is out of our control, since it cannot be handled via standard software.
Table 1: The $\chi^{2}_{\nu}$ tests with 40 bins when comparing the experimental data applying the SNR procedure with simulation. For the ADC distributions with a cut of 100 ADC, and for the distributions related to the number of clusters (i.e., size, and maximum ADC signal) with CC procedure for Sr90 and Cs137 at 0 mm distance between sensor and source for 500 s exposure time.
@@
ADC number (cut >100 ADC) 0.45 0.48
Cluster Size 1.77 1.26
Maximum cluster ADC 2.62 1.94
@@
Figure 11: Cluster size versus maximum ADC signal per cluster for the experimental and simulated data for Sr90 (left) and Cs 137 (right), the experimental data applying the SNR procedure and the CC procedure was applied to experimental data and simulation to compare both, at 0 mm between sensor and source for 500 s exposure time.
In figure 11, we can observe that the simulated data for these distributions are close to the experimental data for both Sr90 and Cs137. Due to the background noise, the experimental data show more scattered clusters above 30 pixels than in the simulation. The $\chi^{2}_{\nu}$ for these distributions is 0.66 for Sr90 and 0.44 for Cs137, showing a good agreement.
In addition, to verify the inverse square distance law behavior of the Sr90 radioactive source we performed measurements every 2 mm from 0 to 18 mm between the source-detector distance, taking 100 frames per each distance selected. In figure 12 we show the average number of clusters at different distances ($z$) for both the experimental data with noise reduction SNR procedure and the simulation and their respective fits to $a/(z+b)^{2}$ (i.e. the inverse square of the distance, where $a$ and $b$ are free parameters)
Figure 12: Average cluster number, as a function of distance ($z$) for the Sr90 source. The experimental data applying the SNR procedure is in red and the pure simulation without cuts is in blue. Both data and simulation are fitted to the inverse square function $y=a/(z+b)^{2}$.
@@
To show the expected behavior of the signal with distance, we apply the SNR procedure to the experimental data, while for both experimental and simulation data the CC procedure was applied, to remove the background. We normalize both distributions to better appreciate their behavior with the source-detector distance as the cuts affect the average cluster count. Figure 13 shows that these adjustments follow the inverse square of the distance. The parameters obtained for these fits are $a=0.16\pm 0.02$, $b=0.41\pm 0.03$ for the experimental data, and for the simulation $a=0.34\pm 0.03$, $b=0.61\pm 0.02$, with a $\chi^{2}_{\nu}$ of 0.018 and 0.014 for the fit of experimental and simulated data, respectively. This demonstrates the expected behavior for the source-detector distance for both experimental data and the simulation. Comparing both fits we get a $\chi^{2}_{\nu}$ of 0.98 this shows that both fits represent the same behavior with the distance.
@@
Figure 13: Average cluster number, as a function of distance ($z$) for the Sr90 source. The SNR procedure was applied to the experimental data, while for both experimental and simulation data the CC procedure was applied. The experimental data is in red and the simulation data is in blue. Both data are fitted to the inverse square function $y=a/(z+b)^{2}$.
To explore the energy resolution capabilities of the commercial CMOS sensor, we used Geant4 simulations to analyze the correlation between the primary electron energy at the point of interaction with the silicon pixel and three parameters: (a) the total energy deposited in the cluster, (b) the maximum energy deposited in a single pixel within a cluster, and (c) the cluster size. The resulting correlations for both radioactive sources are presented in figure 14. No parameter exhibits a clear correlation with the primary electron energy. Within the energy range sensitive to the sensor, delimited by the horizontal black dashed lines, we observe an almost flat response in both energy deposition related parameters with respect to the incident electron’s kinetic energy. This behaviour is even stronger for the cluster size, where small clusters can be produced by both low-energy and high-energy electrons. Interestingly, for very low deposited energies, a slight anticorrelation trend with the primary electron’s kinetic energy is observed. However, this region is shadowed by background contributions and is therefore not accessible for a reliable analysis.
Figure 14: Geant4 simulation of the true kinetic energy of arriving electrons at the silicon pixels emitted by Sr90 (left) and Cs137 (right) versus: a) the total energy deposited in the cluster of pixels, b) the maximum energy deposited in the cluster c) the size of the cluster. The dashed lines are located in the energy corresponding to the ADC lower cut and saturation.
Even if the primary energy cannot be estimated, the sensor setup is capable of accurately describing the energy deposited in the pixel matrix in terms of ADC counts. As seen in figure 7 there was a good agreement between simulation and experimental data. To study this agreement in detail, we compared in figure 15 for Sr90 and Cs137 the counts in bins of 1 ADC, greater than 100 ADC, between experimental data and simulation, normalized by the standard deviation. After fitting both distributions with a Gaussian function, we observe that the errors are almost symmetrical centered around zero, with a sigma of less than one standard deviation.
Figure 15: Distributions of the differences in counts per ADC bin (i.e. deposited energy above 100 ADC) between experimental data and simulation, expressed in units of standard deviations (sigmas), for both Sr90 and Cs137. Each distribution is fitted with a Gaussian function. The fit parameters are shown in the legend.
@@
Given the poor energy resolution of the sensor and following what was seen in figure 8, the sensor has no radioactive source separation power. However, to further test whether the Sr90 and Cs137 sources can be distinguished, we compared the distributions discussed above at the 0-mm distance. Then, we normalize the histograms to avoid differences in the activities of the sources. Figure 16, 17, and 18 show the normalized distributions at the pixel and cluster level after applying the cuts. The $\chi^{2}_{\nu}$ for the comparison between the two sources in the experiment and the simulation is shown in Table 2. The average $\chi^{2}_{\nu}$ for the experimental data is 1.33, and for the simulation, it is 0.25. Thus, it is impossible to distinguish between these two radioactive sources for the camera capabilities, even if the emitted decay radiation spectrum are different.
Table 2: The $\chi^{2}_{\nu}$ test with 40 bins when comparing the experimental data applying the SNR procedure, and simulation of Sr90 with Cs137 for the normalized ADC distributions with a cut of 100 ADC and the normalized distributions related to the number of clusters (i.e. size, and maximum ADC signal) with the CC procedure at 0 mm distance between source and detector for 500 s exposure time.
@@
ADC number (cut >100 ADC) 1.72 0.03
Cluster Size 1.39 0.66
Maximum cluster ADC 0.89 0.06
@@
Figure 17: Normalized cluster size distribution comparing the sources of Sr90 and Cs137 for the experimental data (left) and the simulation (right), experimental data applying the SNR procedure comparing with simulation, and applying the CC procedure to both at 0mm between sensor and source for 500 s exposure time.
Figure 18: Normalized maximum ADC signal per cluster distribution comparing the sources of Sr90 and Cs137 for the experimental data (left) and the simulation (right), experimental data applying the SNR procedure comparing with simulation, and applying the CC procedure to both at 0mm between sensor and source for 500 s exposure time.